OpenStudy (anonymous):

are the complex fractions x^2-x-20/4/x-5/10 and x^2-x-20/x-5/4/10 equivalent? Simplify each, and then explain why or why not.

OpenStudy (danjs):

use parenthesis in your expressions, the way you wrote it is probably not what you meant

OpenStudy (danjs):

(x^2-x-20) /

OpenStudy (danjs):

ok

OpenStudy (danjs):

When there is a fraction divided by another fraction, you can flip the bottom fraction over and multiply instead .. like this...

OpenStudy (danjs):

\[\frac{ x^2 - x - 20 }{ 4 } * \frac{ 10 }{ x - 5 }\]

OpenStudy (danjs):

can you factor the polynomial in the numerator? ...(x ) (x )

OpenStudy (danjs):

factors of 20 that multiply to -20 and add up to -1

OpenStudy (danjs):

\[\frac{ 10 * (x -5)(x+4) }{4 * (x-5) }\]

OpenStudy (danjs):

--Cancel the (x-5) from both numerator and denominator (x-5)/(x-5) = 1 -- Reduce the 10/4 to 5/2

OpenStudy (danjs):

\[\frac{ 5(x+4) }{ 2 }\]

OpenStudy (danjs):

Now do the same thing to the second one..

OpenStudy (mathstudent55):

A complex fraction can be simplified this way: \(\Large \dfrac{\frac{\color{red}{a}}{\color{green}{b}}}{\frac{\color{green}{c}}{\color{red}{d}}} = \dfrac{\color{red}{ad}}{\color{green}{bc}}\) \(\Large \dfrac{\frac{x^2 - x - 20}{4} }{\frac{x-5}{10} } = \dfrac{10(x^2 - x - 20)}{4(x - 5)}\) \(\Large \dfrac{\frac{x^2 - x - 20}{x - 5} }{\frac{4}{10} } = \dfrac{10(x^2 - x - 20)}{4(x - 5)}\) The complex fractions are equivalent.

OpenStudy (anonymous):

Thank you both! :)

OpenStudy (mathstudent55):

You're welcome.